OpenStudy (anonymous):

Which equation is equivalent to 4(3 - 5x) = 7 - 2(1-4x)? A. 8x = 5 B. 8x = 17 C. 29x = 5 D. 29x = 17 E. None of the above

OpenStudy (anonymous):

none of the above correct

OpenStudy (jiteshmeghwal9):

4(3-5x)=7-2(1-4x) 12-20x=7-2+8x 12-20x=5+8x now take variables at one side & constants at one side as -20x-8x=5-12 -28x=-7 x=-7/-28=1/4. therefore,none of the above.